# Cross-Problem Parameter Transfer in Quantum Approximate Optimization Algorithm: A Machine Learning Approach

### v1

@@ Abstract
Reproducibility: Our results and code are available at [link will be added upon acceptance]
@@
Figure 1: Overview of the parameter transfer procedure. First, QAOA is run on a quantum circuit to optimize for a donor COP (MaxCut) and retrieve the optimal parameter set $(\boldsymbol{\gamma}^{*},\boldsymbol{\beta}^{*})$. Next, those parameters are directly applied to the acceptor COP (MIS) without additional tuning.
@@ III-A Quantum Approximate Optimization Algorithm
where $\text{Pr}(x)$ is the probability of observing $x$ when measuring all qubits of |$\ket{\boldsymbol{\gamma},\boldsymbol{\beta}}$. The approximation ratio is then calculated as
@@ III-C Graph Embedding
In this paper, we design a retrieval model that learns whole-graph embeddings to identify suitable donors for parameter transfer, taking inspiration from information retrieval, such as text-based search [33] and recommendation systems [34]. Specifically, we implement a GNN that predicts a transfer score for any donor–acceptor pair, and during inference, we use these predicted scores to retrieve the donor with the highest potential for effective parameter transfer.
@@ V-C Results on Parameter Transferability
In this experiment, we evaluate the quality of the donor graphs identified by our machine learning model. Once the top-5 donors are chosen for each acceptor graph in the test set $\mathcal{S}_{\text{test}}$, we take each donor’s 16 optimized MaxCut parameters, apply them to the MIS circuit, and perform QAOA inference for 1,000 shots to measure the approximation ratio $r$ and the probability of finding the optimal solution $\text{probability}(^{*}x)$. We record the best-performing parameter set among the five donors based on the metrics and present the results in Tab. I.
Regarding the mean best approximation ratio $r$ for $p=2$, ChebConv and Graph2Vec achieve the best results with 0.8370 and 0.8385. The Closeness approach delivers fair results with an approximation ratio around 0.80, but generally is still the lowest among all methods, highlighting the importance of further fine-tuning the original Graph2Vec embeddings to harness the full potential. We further illustrate the results for ChebConv by graph types and number of nodes in Fig. 3. We see that Barabási–Albert graphs achieve the lowest overall approximation ratio, and the mean best approximation ratio gradually decreases as the node count grows, signifying the increased difficulty of transferring parameters to larger graphs. A similar pattern is observed in the probability of finding the optimal solution, indicating that Barabási–Albert to be a particularly challenging class of graph for parameter transfer from MaxCut to MIS.
Figure 5: For $p=2$, using the best transferred MaxCut parameters with $r\geq 0.80$ as initializations for MIS, we further optimize the QAOA circuit for just 10 additional steps. This brief post-optimization phase boosts the average approximation ratio from 0.8933 to 0.9500.
@@ V-D Results on Optimization Speedup
In this subsection, we show how leveraging the transferred MaxCut parameters can significantly accelerate the optimization process and yield higher approximation ratios compared to starting from random initializations. We first select the best transferred parameters that yield an approximation ratio $r\geq 0.8$ and use them as intializations for the MIS circuit. We then perform QAOA optimization for only 10 steps. This brief post-optimization phase significantly boosts the average approximation ratio to 0.9500, as shown in Fig. 5. For comparison, we also optimize each MIS circuit from 16 random initializations and record that it takes 155 steps to reach an average approximation ratio of 0.8833. These results underscore the gains in efficiency and performance of parameter transfer, demonstrating how reusing carefully chosen MaxCut parameters substantially cuts down on optimization time while producing superior MIS solutions compared to random initializations.
@@ VI Conclusion and Future Outlook
Reproducibility: Our results and code are available at [link will be added upon acceptance]

### v2

@@ Abstract
Reproducibility: Our results and code are available at this link: https://github.com/Nyquixt/Cross-Problem-PT-QAOA.
@@
Figure 1: Overview of the parameter transfer procedure. First, QAOA is run on a quantum circuit to optimize for a donor COP (i.e. MaxCut) and retrieve the optimal parameter set $(\boldsymbol{\gamma}^{*},\boldsymbol{\beta}^{*})$. Next, those parameters are directly applied to the acceptor COP (i.e. MIS) without additional tuning.
@@ III-A Quantum Approximate Optimization Algorithm
where $\text{Pr}(x)$ is the probability of observing $x$ when measuring all qubits of |$\ket{\boldsymbol{\gamma},\boldsymbol{\beta}}$. The approximation ratio is calculated as
@@ III-C Graph Embedding
In this paper, we design a retrieval model that learns whole-graph embeddings to identify suitable donors for parameter transfer, taking inspiration from information retrieval, such as text-based search [33] and recommendation systems [34]. Specifically, we implement a neural network that predicts a transfer score for any donor–acceptor pair, and during inference, we use these predicted scores to retrieve the donor with the highest potential for effective parameter transfer.
@@ V-C Results on Parameter Transferability
We evaluate the quality of the donor graphs identified by our machine learning model. We pick $k=$ 5 following the standard practice of the instance retrieval research in machine learning [47]. Once the top-5 donors are chosen for each acceptor graph in the test set $\mathcal{S}_{\text{test}}$, we take each donor’s 16 optimized MaxCut parameters, apply them to the MIS circuit, and perform QAOA inference for 1,000 shots to measure the approximation ratio $r$ and the probability of finding the optimal solution $\text{probability}(^{*}x)$. We record the best-performing parameter set among the five donors based on the metrics and present the results in Tab. I.
Regarding the mean best approximation ratio $r$ for $p=2$, ChebConv and Graph2Vec achieve the best results with 0.8370 and 0.8385. The Closeness approach delivers fair results but generally is still the lowest among all methods, highlighting the importance of further fine-tuning the original Graph2Vec embeddings to harness the full potential. We further illustrate the results for ChebConv by graph types and number of nodes in Fig. 3. We see that Barabási–Albert graphs achieve the lowest overall approximation ratio, and the mean best approximation ratio gradually decreases as the node count grows, signifying the increased difficulty of transferring parameters to larger graphs. A similar pattern is observed in the probability of finding the optimal solution, indicating that Barabási–Albert to be a particularly challenging class of graph for this setting.
Figure 5: For $p=2$, using the best transferred MaxCut parameters with $r\geq$ 0.8 as initializations for MIS, we further optimize the QAOA circuit for just 10 additional steps. This brief post-optimization boosts the average approximation ratio from 0.8933 to 0.9500. The results are from the ChebConv-based model.
@@ V-D Results on Optimization Speedup
In this subsection, we show how leveraging the transferred MaxCut parameters can significantly accelerate the optimization process and yield higher approximation ratios compared to starting from random initializations. We first select the best transferred parameters that yield an approximation ratio $r\geq$ 0.8, use them as intializations for the MIS circuit, and perform QAOA optimization for only 10 steps. This brief post-optimization phase significantly boosts the average approximation ratio to 0.9500. In addition, we plot our result against the average approximation ratio of direct transfer of MaxCut parameters without further optimization, random initialization of MIS parameters and random initialization of MIS parameters with 10-step optimization in Fig. 5. For comparison, we also optimize each MIS circuit from 16 random initializations and record that it takes 155 steps to reach an average approximation ratio of 0.8833. These results underscore the gains in efficiency and performance of parameter transfer, demonstrating how reusing carefully chosen MaxCut parameters substantially cuts down on optimization time while producing superior MIS solutions compared to random initializations.
@@ VI Conclusion and Future Outlook
Reproducibility: Our results and code are available at this link: https://github.com/Nyquixt/Cross-Problem-PT-QAOA.